OpenStudy (anonymous):

can someone help me differentiate the dx at the end of a bracket of this equation...

OpenStudy (anonymous):

(1-6/x^3)dx

OpenStudy (anonymous):

Are you trying to differentiate the whole term?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Ok. d/dx 1 - 6 d/dx 1/x^3 = -6 * -3/x^4 =18/x^4

OpenStudy (anonymous):

ok so one is nothing right?

OpenStudy (anonymous):

Correct. d/dx 1 is 0

OpenStudy (anonymous):

thank you ill give you a medal :)

OpenStudy (anonymous):

You're very welcome.